# Timestamp Manipulation: Timestamp-based Nakamoto-style Blockchains are Vulnerable

### v4

@@ Abstract
Nakamoto consensus are the most widely adopted decentralized consensus mechanism in cryptocurrency systems. Since it was proposed in 2008, many studies have focused on analyzing its security. Most of them focus on maximizing the profit of the adversary. Examples include the selfish mining attack [FC ’14] and the recent riskless uncle maker (RUM) attack [CCS ’23]. In this work, we introduce the Staircase-Unrestricted Uncle Maker (SUUM), the first block withholding attack targeting the timestamp-based Nakamoto-style blockchain. Through block withholding, timestamp manipulation, and difficulty risk control, SUUM adversaries are capable of launching persistent attacks with zero cost and minimal difficulty risk characteristics, indefinitely exploiting rewards from honest participants. This creates a self-reinforcing cycle that threatens the security of blockchains. We conduct a comprehensive and systematic evaluation of SUUM, including the attack conditions, its impact on blockchains, and the difficulty risks. Finally, we further discuss four feasible mitigation measures against SUUM.
publicationid: pubid: Network and Distributed System Security (NDSS) Symposium 2026 23-27 February 2026, San Diego, CA, USA ISBN 979-8-9894372-8-3 https://dx.doi.org/10.14722/ndss.2026.250025 www.ndss-symposium.org
@@ I Introduction
The Proof-of-Work (PoW) blockchains [1, 2], with Bitcoin [3] and Ethereum 1.x [4, 5] as prominent examples, have significantly propelled the thriving development of the cryptocurrency sector, which boasts a market valuation of $1.9$ trillion dollars [6]. These advanced blockchain systems are built upon highly decentralized blockchain protocols and ingeniously employ the PoW mechanism to ensure their consistency and security [7, 8, 9, 10, 11, 12], and the consensus mechanism is also called Nakamoto consensus. The blockchain network is jointly maintained and operated by a group of active participants known as miners, who possess and control a certain amount of computational resources, such as high-performance computers and specialized mining hardware. Participants invest substantial computational power to solve elaborately designed and extremely complex cryptographic puzzles, a process referred to as mining [13]. Participants who successfully solve these puzzles gain the right to record the next valid block on the blockchain.
The blockchain system provides incentives to participants who successfully generate valid blocks, to ensure the security of the consensus mechanism. These incentives typically consist of a certain quantity of newly issued cryptocurrencies and transaction fees included in the block. Ideally, the reward allocation mechanism designed by the blockchain system is relatively equitable, accurately reflecting the proportion of computational resources invested by participants [14, 15, 16]. This implies that participants with more computational resources will have a higher probability of mining blocks and receiving corresponding coin-base rewards, thereby incentivizing them to continue contributing to the stable operation and security of the blockchain network.
@@
Ethereum PoW [17], Ethereum Fair [18] and Ethereum Classic [19] have evidently gained popularity among participants: as of April 1, 2025, their combined hash power accounted for $13.9\%$ of the peak hash power observed in the Ethereum 1.x ecosystem [20]. Ethereum 1.x boasts a significantly reduced block time of approximately $13$ seconds on average, compared to Bitcoin’s average block time of 10 minutes, which results in frequent state fork (also known as soft fork) [21]. A series of mining attacks aimed at maliciously manipulating block difficulty have been proposed and proven effective in reducing blockchain security [22, 23, 24, 25, 16, 26, 27]. Among them, the most recent one is the riskless Uncle Maker attack (RUM). The idea is to manipulate the timestamp of a proposed block to make its difficulty higher than that of honest blocks in other branches, so as to achieve the advantage of higher profits. The attack is risk-free where it incurs no additional overhead. By examining Ethereum 1.x blockchain data, the authors of RUM provided compelling evidence indicating that participants executed variants of the RUM attack for approximately two years [28]. One of the co-founders of F2Pool (then the second-largest mining pool for Ethereum 1.x) acknowledged their manipulation of block timestamps in this manner, marking the first instance of consensus protocol manipulation by participants in practice [29].
Our work further extend RUM by introducing two advanced attack strategies, the Unrestricted Uncle Maker (UUM) Attack and the Staircase-Unrestricted Uncle Maker (SUUM) Attack, that fundamentally threaten the security of timestamp-based Nakamoto-style blockchains by inflicting permanent systemic harm. Unlike prior work that merely enhances adversarial rewards, these attacks exploit vulnerabilities in timestamp manipulation and fork selection rules to irreversibly destabilize blockchain fairness and incentive mechanisms. Specifically, The core design of UUM and SUUM also exploit vulnerabilities in fork selection rules to turn honest blocks in other forks into uncle blocks. Unlike RUM, UUM abandons the zero-cost constraint, expanding the scope of attack initiation conditions to achieve higher returns than RUM while incurring nearly no costs. As an upgraded version of UUM, SUUM strategically combines block withholding/releasing with malicious timestamp manipulation to enable zero-cost, minimal-risk persistent attacks, ultimately forming a self-reinforcing cycle: they suppress honest participants’ returns by manipulating difficulty growth rates while amplifying their own profits through persistent chain reorganizations. To quantify UUM/SUUM’s dominance, we implement a discrete-event timestamp-based Nakamoto-style blockchan simulator and conduct large-scale experiments (1M blocks, 10K trials). Through it, We validate that SUUM adversaries achieve disproportionate reward advantages (e.g., 33.30% vs 28.41% vs 26.12% for SUUM, UUM and RUM at adversary’s relative power $\alpha=0.25$) over both honest mining and prior attacks, all while maintaining minimal difficulty risks through precise timestamp calibration. It is noteworthy that our proposed attacks are applicable to any blockchain system that adopts the Ethereum 1.x’s fork selection rule, including Ethereum PoW [17], Ethereum Fair [18], Ethereum Classic [19], and others.
The cumulative effect of these strategies permanently erodes the protocol’s security assumptions. We analyze the attack using a zero-sum game as a game between UUM/SUUM adversary and honest participants. Our simulations demonstrate that even adversaries with modest computational power (e.g., $\alpha=0.3$) reduce honest participants’ rewards by 11.85% compared to baseline honest mining. This imbalance escalates exponentially as adversarial power grows, ultimately leading to the collapse of the protocol’s economic model, where rational participants are incentivized to join adversarial coalitions rather than mine honestly.
@@ 4. $\bullet$
    Systemic Economic Collapse. We prove SUUM triggers a death spiral: adversarial rewards scale super-linearly, incentivizing rational miners to defect. This accelerates protocol abandonment, ultimately collapsing the economic model.
@@ 1. $\bullet$
To facilitate a deeper understanding of the composition of rewards, we employ an illustrative example to delineate the distinctions among various types of rewards, as depicted in Figure 2.
Example 2 (Block Types).
Solid white blocks signify regular blocks that constitute the longest valid chain; their creators are entitled to corresponding coin-base rewards. Solid gray blocks represent uncle blocks, which are the initial blocks linking a regular block on a non-longest valid chain. When an uncle block is referenced by a subsequent regular block, its miners are eligible for uncle rewards. Solid blue blocks encompass both regular and nephew blocks, situated on the longest valid chain and referencing other uncle blocks; their miners receive coin-base rewards and nephew rewards. Conversely, dashed white blocks symbolize orphan blocks, for which their creators receive no rewards, serving as a penalty for deviant mining practices.
@@ VI Comparison
This result once again implies that adversarys can relatively easily carry out UUM and SUUM attacks with almost no cost risk, thereby further highlighting the severe threats posed by these two attack strategies to the security of timestamp-based Nakamoto-style blockchains. Since nearly cost-free attacks may prompt more adversarys to attempt to exploit these vulnerabilities, undermining the fairness and stability of the blockchain.
@@ VIII Discussion
We now analyze the essence of the Uncle Maker attack and propose some mitigation measures to effectively prevent the Uncle Maker attack and its advanced variants. These measures mainly focus on adjusting the consensus mechanism and incentive mechanism of the blockchain, aiming to reduce the profit space of adversaries and enhance the security and fairness of the timestamp-based Nakamoto-style blockchains. The details can be found in Appendix N

### v5

@@ Abstract
Nakamoto consensuses are the most widely adopted decentralized consensus mechanism in cryptocurrency systems. Since it was proposed in 2008, many studies have focused on analyzing its security. Most of them focus on maximizing the profit of the adversary. Examples include the selfish mining attack and the recent riskless uncle maker (RUM) attack. In this work, we introduce the Staircase-Unrestricted Uncle Maker (SUUM), the first block withholding attack targeting the timestamp-based Nakamoto-style blockchain. Through block withholding, timestamp manipulation, and difficulty risk control, SUUM adversaries are capable of launching persistent attacks with zero cost and minimal difficulty risk characteristics, indefinitely exploiting rewards from honest participants. This creates a self-reinforcing cycle that threatens the security of blockchains. We conduct a comprehensive and systematic evaluation of SUUM, including the attack conditions, its impact on blockchains, and the difficulty risks. We conducted an empirical analysis on three major ETH 1.x-based blockchains, namely Ethereum 1.x, Ethereum Classic, and Ethereum PoW. Statistical results suggest that, as of October 2025, four major mining pools had been launching SUUM attacks by manipulating timestamps. Finally, we further discuss four feasible mitigation measures against SUUM.
@@ I Introduction
Timestamp-based Nakamoto-style Blockchains. Nakamoto-style blockchains, with Bitcoin [1] and Ethereum 1.x [2, 3] as prime examples, have catalyzed the thriving cryptocurrency sector [4]. These systems rely on consensus mechanisms like Proof-of-Work (PoW) to ensure security and consistency [5, 6, 7, 8, 9, 10]. Miners invest computational resources to solve cryptographic puzzles and append new blocks to the chain [11]. The protocol’s incentive mechanism is designed to be fair, rewarding participants in proportion to their contributed computational power [12, 13, 14]. Ethereum 1.x and its popular derivatives (Ethereum PoW [15], Ethereum Fair [16], Ethereum Classic [17]) feature a significantly reduced average block time of 13 seconds compared to Bitcoin’s 10 minutes. This higher throughput, however, leads to frequent soft forks [18], creating opportunities for mining attacks that maliciously manipulate block difficulty to gain an unfair advantage [19, 20, 21, 22, 23, 24, 14].
Existing Attack Models. A significant body of research has focused on identifying and analyzing vulnerabilities in blockchain incentive models. A prominent line of work explores attacks such as selfish mining [25] and its variants. Recently, the Riskless Uncle Maker (RUM) attack [26] demonstrated a novel exploitation vector: timestamp manipulation. Unlike selfish mining that relies on block withholding alone, RUM exploits timestamp rules to inflate block difficulty, achieving risk-free profits. RUM adversaries strategically set block timestamps to artificially inflate their block’s difficulty above that of competing honest blocks, thereby ensuring preferential selection during forks. This attack is proven to be risk-free, incurring no additional overhead.
Limitations and Challenges. The RUM attack constitutes a tangible and impactful threat [26]. Despite this, its threat model is fundamentally constrained by a rigid, risk-averse design. This conservative posture is not a minor oversight but a structural limitation. It directly restricts the attack’s scope and long-term impact, thereby failing to reflect the full spectrum of adversarial behavior in decentralized networks.
Compounding this issue are gaps in how prior work models attack synergy. The existing security landscape thus contains two critical, unaddressed vulnerabilities. Both of these vulnerabilities are directly tied to the practical risks and adversarial incentives detailed in this paper’s subsequent sections:
L1: Constrained Adversarial Capability.
The RUM attack has a core requirement of zero additional difficulty risk [26]. This requirement forces it to rely on an extremely narrow set of attack initiation conditions. These conditions rarely align with the actual block generation dynamics of Ethereum 1.x (we will quantify this later in the simulations in Section VII). Specifically, the RUM attack only triggers under one condition. The condition is the timestamp difference between a newly mined honest block ($B_{1}^{p_{h}}$) and its parent mainchain block ($B_{0}^{p_{h}}$) must fall within the [9, 18) second window. Mathematically, this corresponds to $\left\lfloor\frac{t_{1}^{p_{h}}-t_{0}^{p_{h}}}{9}\right\rfloor=1$. This constraint is not arbitrary. It ensures the RUM attack avoids any upward fluctuation in the blockchain’s long-term difficulty, thus maintaining its zero-risk feature. Yet this rigidity directly limits the RUM attack’s impact. We will demonstrate this limitation in detail in our subsequent empirical analysis: $\bullet$ Low attack frequency. Ethereum 1.x’s average block time is 13 seconds [2], meaning the [9, 18) second timestamp gap (required for RUM to initiate) occurs in only 12% of honest block pairs (Section VII-A’s steady-state probability calculations confirm this). This caps RUM’s steady-state attack probability at just 1.74% (vs. SUUM’s 16.84%), severely limiting its reward extraction potential. $\bullet$ Capped profitability. RUM’s risk-free constraint prevents it from exploiting scenarios where even minimal difficulty risk could double its attack opportunities. As we later show in Theorem 5, this leaves RUM’s maximum reward share at 26.12% (for the adversary’s relative power $\alpha=0.25$), far below the 33.30% achieved by risk-tolerant strategies like SUUM.
This mismatch between RUM’s assumptions and real-world adversarial incentives. Miners always tolerate tiny, controllable risks for higher returns, which means existing threat models underestimate how adversaries could scale timestamp-based attacks by relaxing risk constraints:
L2: Narrow Attack Vector Integration.
All prior analyses of timestamp manipulation, including those on the RUM attack, treat this technique in isolation. They completely overlook its potential synergy with block withholding. Block withholding is a well-documented strategy in selfish mining [25] and stubborn mining [27]. However, it has never been combined with timestamp exploitation. Such isolated analysis leads to a severe underestimation of attack impact. We will later prove this through the SUUM attack (Section V) and validate it via simulations (Section VII-D): $\bullet$ Single vs. cascading harm. The RUM attack uses timestamp manipulation alone. Each attack can only replace one honest block, turning it into an uncle block. However, when combined with block withholding, the situation changes. Adversaries first build a private chain using withheld blocks. They then release these blocks. In this case, timestamp manipulation ensures every block in the private chain has higher difficulty than honest blocks at the same height. As shown in Figure 3-(c), this results in cascading chain reorganizations. A single SUUM attack can invalidate multiple honest blocks compared to only one block invalidated by RUM. When the adversary’s relative power $\alpha=0.3$, honest rewards decrease by 11.85% (Section VII-B). $\bullet$ Loss of persistence. The RUM attack relies on an isolated strategy. It cannot sustain attacks over the long term. The reason is that it lacks a mechanism to lock in advantages. By contrast, the SUUM attack uses block withholding to create a self-reinforcing cycle. Withheld blocks with manipulated timestamps suppress the growth of honest blocks’ difficulty (Section V-C). At the same time, chain reorganizations allow the adversary to steal rewards. This makes it economically rational for honest miners to join the attack. We formalize this death spiral in Theorem 8.
Failure to model such synergy leads to a critical issue. Current threat models cannot capture the systemic and permanent harm. This harm is inflicted on the blockchain’s economic model by coordinated strategies. We address this gap through SUUM’s integrated framework. The above limitations leave fundamental questions unresolved:
Q1: Can RUM adversaries systematically relax the constraints to achieve higher rewards and greater persistence?
Q2: Does there exist a synergistic attack strategy that combines timestamp manipulation with block withholding to inflict permanent, systemic harm on the protocol’s economic model?
Proposed Attack Frameworks. To address these questions, we propose two advanced attack strategies that fundamentally extend the threat model for timestamp-based Nakamoto-style blockchains:
S1 to Q1: Unrestricted Uncle Maker (UUM) Attack.
The UUM attack strategically abandons the zero-cost constraint of RUM, expanding the scope of attack initiation conditions to achieve higher returns. By tolerating minimal, carefully calibrated difficulty risk, UUM adversaries significantly increase their steady-state probability of being in an attack state, thereby amplifying their reward share compared to RUM.
S2 to Q2: Staircase-Unrestricted Uncle Maker (SUUM) Attack.
As a strategic escalation of UUM, the SUUM attack integrates block withholding and release mechanisms with precision timestamp manipulation. This synergy enables persistent, near zero-cost attacks. SUUM forms a self-reinforcing cycle: it suppresses honest participants’ returns by manipulating difficulty growth rates while amplifying adversarial profits through strategic chain reorganizations. We formalize SUUM via a three-pillar framework: 1) timestamp manipulation for difficulty dominance, 2) block withholding for reorganization cascades, and 3) difficulty risk control via temporal gap optimization.
UUM Attack. The UUM attack operationalizes a three-phase framework to achieve low-risk, high-reward objectives by capitalizing on timestamp manipulation within Ethereum 1.x-style fork selection rules. In the Deployment Phase, the adversary monitors the timestamp difference between a newly generated honest block and its preceding mainchain block, waiting for this difference to be greater than or equal to 9 seconds. This relaxed condition, compared to RUM’s narrower window, significantly increases the frequency of attack opportunities. Once the condition is met, the system enters the Execution Phase. Here, the adversary mines on the parent block and, upon finding a block before honest participants, sets its timestamp to create a difficulty advantage. Specifically, the adversary ensures the timestamp difference for its block is between 1 and 900 seconds, making its difficulty higher than the competing honest block and triggering a successful fork.
In the Minimum Risk Control Phase, the adversary meticulously calibrates the timestamp to maximize the difficulty advantage while minimizing the subsequent impact on the blockchain’s long-term difficulty growth, often by setting the floor function of the timestamp difference to 1. This phased approach allows UUM to trade a minimal, controlled amount of risk for a substantially higher steady-state attack probability and reward yield than RUM.
@@
SUUM Attack. Building on UUM’s risk-tolerant design, SUUM further integrates block withholding to enable persistent, cost-free attacks. It is orchestrated through a sophisticated five-phase model. The attack begins in the SUUM Deployment Phase. If the adversary finds a block first, it withholds it, transitioning to the SUUM Block Withholding Phase and building a private chain. This mirrors selfish mining [25] but is augmented with precise timestamp control. If an honest miner finds a block that meets certain timestamp conditions (timestamp difference $\geq$ 9 seconds), SUUM can Downgrade to a UUM-like strategy, ensuring it remains profitable even when not leading.
The core of SUUM’s power lies in the SUUM Releasing Block Phase. When an honest block is published, the adversary immediately releases one or more withheld blocks from its private chain. Crucially, each released block has its timestamp manipulated to ensure its difficulty exceeds that of the honest chain at the same height, causing a chain reorganization that turns honest blocks into uncles or orphans. This process is governed by the SUUM Minimum Risk Control Phase, where the adversary sets the timestamps of its withheld blocks to be just high enough to win fork competitions while minimizing the resultant difficulty increase, thus preserving its future mining efficiency. This combination of withholding and timestamp manipulation creates a staircase effect, where the adversary can repeatedly release blocks to reorganize the chain, maximizing rewards and inflicting sustained damage on honest miners’ profitability.
Our Results. Through extensive discrete-event simulations of a timestamp-based Nakamoto-style blockchain, we empirically validate that SUUM achieves superior reward extraction (e.g., 33.30% vs. 28.41% for UUM and 26.12% for RUM at adversary’s relative power $\alpha=0.25$) and induces a higher forking rate (16.84% vs. 13.40% for UUM and 1.74% for RUM). Furthermore, we demonstrate that SUUM triggers a death spiral in the protocol’s economic model, where rational miners are incentivized to join adversarial coalitions, accelerating protocol abandonment.
Latest empirical evidence from mainstream ETH 1.x-style blockchains further underscores the urgency of timestamp manipulation threats. We analyzed 50,130 mainchain blocks across three prominent networks: Ethereum 1.x (block heights 15505647–15535776), Ethereum Classic (23232147–23242146), and Ethereum PoW (22905613–22915612). We identified the presence of persistent, systematic timestamp manipulation.
Key observations from real-world data across all three blockchains include: 1) The proportion of blocks with timestamp differences divisible by 9 is far lower than the expected normal ratio; 2) The proportion of blocks with timestamp differences equal to multiples of 9 minus 1 is significantly higher than the expected normal ratio; 3) Four major mining pools (Ethereum 1.x: 0x829bd8…, Ethereum PoW: 0x9205c2…, Ethereum Classic: 0x406177… and 0x35aa26…) perfectly avoid generating blocks with timestamp differences that are multiples of 9. This deliberate timestamp manipulation aims to increase the probability of winning in fork competitions and avoid block difficulty fluctuations, which highly aligns with the minimal risk control strategy (Theorem 7) in our proposed SUUM/SUM attacks. These findings confirm that timestamp manipulation is not a transient issue limited to the history of Ethereum 1.x, but a persistent and widespread threat across all ETH 1.x-style blockchains.
@@ 5. $\bullet$
    Exposing Timestamp Manipulation Attack Pools. Through an investigation of three major Ethereum 1.x-style blockchains, we confirm that timestamp manipulation attacks have been implemented in real-world environments, and expose four mining pools that are currently executing such attacks.
@@ III-A Preliminaries
Definition 1 formally defines the fairness of blockchains. Over the long term, in accordance with the Law of Large Numbers, the reward distribution among participants strictly converges to their resource input (power/stake). This definition not only accounts for the characteristics of blockchain, such as intermittent block generation and resource fluctuations, but also eliminates the interference of short-term random fluctuations via the limit operation. It serves as a core quantitative criterion for evaluating the rationality of incentive mechanisms and whether attacks undermine system fairness.
@@ VI Comparison
This result once again implies that adversaries can relatively easily carry out UUM and SUUM attacks with almost no cost risk, thereby further highlighting the severe threats posed by these two attack strategies to the security of timestamp-based Nakamoto-style blockchains. Since nearly cost-free attacks may prompt more adversaries to attempt to exploit these vulnerabilities, undermining the fairness and stability of the blockchain.
@@ VII-D Estimate the Forking Rate
VII-E Empirical Validation on Real Blockchain Networks
Previous sections have validated the theoretical effectiveness of SUUM/UUM attacks through discrete-event simulations (Sections VII.A–VII.D). However, the real-world existence of such attacks requires corroboration from actual blockchain data. This section leverages on-chain data from three ETH 1.x-style blockchains, Ethereum (ETH, block heights 15505647–15535776, 30,130 mainchain blocks), Ethereum Classic (ETC, block heights 23232147–23242146, 10,000 mainchain blocks), and Ethereum PoW (block heights 22905613–22915612, 10,000 mainchain blocks), to address three key questions via two experiments: 1) Whether the distribution of timestamp differences aligns with the attack window characteristics of SUUM/UUM; 2) Whether top mining pools engage in systematic timestamp manipulation; 3) Whether certain pools deliberately avoid timestamp differences divisible by 9. Ultimately, these experiments confirm the real-world pervasiveness of SUUM/UUM attacks.
Timestamp Difference in Mainchain Blocks. We counted the number of mainchain blocks corresponding to timestamp differences from parent blocks across the three blockchains, with visualizations presented in Fig. 6a. The statistical results indicate that the proportion of mainchain blocks with timestamp differences divisible by 9 is significantly lower than the expected normal trend. This corroborates that some mining pools are still maliciously manipulating timestamps to gain disproportionate advantages. Furthermore, we observed that the decrease in the number of blocks with timestamp difference $t$ exactly matches the increase in the number of blocks with difference t1t-1. This pattern perfectly corresponds to the difficulty risk control mechanism discussed in previous sections.
Four Timestamp Manipulation Pools. For each of the three blockchains, we sorted mining pools in descending order of total blocks mined and selected the top 10 pools. Among these 30 top pools, we identified four pools that actually implement timestamp manipulation: Ethereum 1.x (0x829bd8…), Ethereum PoW (0x9205c2…), and Ethereum Classic (0x406177… and 0x35aa26…), with their timestamp difference distributions visualized in Fig. 6b. Notably, none of their timestamp difference distributions include values divisible by 9. When a mainchain block would otherwise have a timestamp difference divisible by 9, these pools maliciously adjust it to the original value minus 1. This minimizes difficulty risk while securing a dominant advantage in fork competitions, aligning perfectly with the proposed SUUM/SUM attack strategies.
(a) Three ETH 1.x Blockchains
(b) Four Malicious Pools
Figure 6: Circumstantial Evidence for Attack. (a) This figure presents the distribution of mainchain blocks across three ETH 1.x-style blockchains, Ethereum 1.x, Ethereum PoW, and Ethereum Classic with respect to their timestamp differences from parent blocks. Blocks corresponding to timestamp differences of 9, 18, and 27 seconds are significantly fewer than those in adjacent intervals, while blocks with differences of 8, 17, and 26 seconds exhibit a marked surge. This serves as direct evidence of mining pools’ deliberate manipulation to avoid timestamp differences divisible by 9. (b) This figure focuses on four malicious mining pools, presenting the distribution of timestamp differences among mainchain blocks mined by these pools. All four pools exhibit zero blocks with timestamp differences of 9, 18, and 27 seconds, while blocks corresponding to differences of 8, 17, and 26 seconds show a significant surge. This distribution characterized by the avoidance of multiples of 9 and concentration on values equal to multiples of 9 minus 1 serves as direct evidence of the strategy in SUUM attacks to minimize difficulty risks and maximize competitive advantages in fork contests. It confirms that these pools are systematically executing timestamp manipulation attacks.
@@ VIII Discussion
We now analyze the essence of the timestamp manipulation and propose some mitigation measures to effectively prevent it and its advanced variants. These measures mainly focus on adjusting the consensus mechanism and incentive mechanism of the blockchain, aiming to reduce the profit space of adversaries and enhance the security and fairness of the timestamp-based Nakamoto-style blockchains. The details can be found in Appendix N